40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 15 12:25:38 2011, kiwamu, Update, Green Locking, Y arm ALS : beat-note free run fluctuation noise_budget.png
    Reply  Wed Nov 16 17:29:46 2011, kiwamu, Update, Green Locking, Some updates on the Y end green PDH Yarm_fluctuation.png
       Reply  Thu Nov 17 17:21:22 2011, kiwamu, Update, Green Locking, Y end green PDH servo : it's okay servoTF.pngservofilter.pngerr_suppression.png
          Reply  Fri Nov 18 11:28:39 2011, rana, Update, Green Locking, Y end green PDH servo : it's okay 
             Reply  Fri Nov 18 12:11:24 2011, Zach, Update, Green Locking, Y end green PDH servo : it's okay 
    Reply  Thu Nov 17 18:20:26 2011, kiwamu, Update, Green Locking, Y arm ALS : 1st trial of noise budget Yarm_ALS_2011Nov16.png
       Reply  Mon Nov 21 16:08:04 2011, kiwamu, Update, Green Locking, 2nd trial of Y arm ALS noise budget : broad band noise gone Yarm_ALS_2011Nov19.png
          Reply  Tue Nov 22 15:18:18 2011, kiwamu, Update, Green Locking, broad band noise depends on the gain of Y green PDH. and comaprator broken 
Message ID: 5929     Entry time: Thu Nov 17 17:21:22 2011     In reply to: 5914     Reply to this: 5945
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Y end green PDH servo : it's okay 

Quote from #5914
So I have added an SR560 in the other input of the Newfocus servo box to make the filter shape 1/f^2.
I will post the servo shape and diagram later.

The Y arm green PDH servo is working fine with a sufficient amount of suppression.

(Servo filters)

 As reported on the previous elog entry (#5914) an SR560 was installed to provide one more pole-zero combination in the servo filter.
Here is a plot showing the transfer function of the latest servo filter.
   servoTF.png

And the servo configuration looks like this :

  servofilter.png

 The demodulated signal is split into two path; one goes directly to the Newfocus servo box and the other goes through SR560.
With the SR560 the two way summing path makes a pole at 1 Hz and zero at 100 Hz with when the SR560 has a gain of 100.
The overall gain is adjustable from a knob on the Newfocus servo box.
 

(the Error signal)

 One of the reasons we wanted to increase the servo gain was that :
the laser frequency has to be tightly locked to the Y arm motion because the laser frequency must represent the arm motion in our scheme.
 
Our requirement for allowing a successful ALS is : RMS < 10 pm (1/100 of the cavity linewidth)

I took a spectrum of the error signal when the laser was locked to the Y arm and found that it meets the requirement.

   err_suppression.png

 In the plot I also put a dark noise from the PD to make sure the in-loop noise is above the dark noise.
Right now the power lines at 60 Hz and 180 Hz are lifting the RMS up.
Note that the UGF was at 20-30 kHz.
ELOG V3.1.3-